OpenStudy (anonymous):

Helppp :( You invest $25/ month at a rate of 3.25% APR compounded monthly for 30 years. I need to find its contributions and interest earned. I already have the final balance: $15,209.31

OpenStudy (anonymous):

Savings/ investment calculator: http://www.mycalculators.com/ca/savecalcm.html

OpenStudy (anonymous):

you can't use we excel?

OpenStudy (anonymous):

I tried, but my computer exits out everytime @a-higbee

OpenStudy (anonymous):

\[2[2 + \log _{2}3]-\log _{2}^{2}+\log _{2}3\]

OpenStudy (anonymous):

I haven't done those type of problems in the longest, so I hardly remember how to solve them or to even start them :/ @a-higbee